OpenStudy (anonymous):

will u integrate this (x+6)dx/(x+2)^2

OpenStudy (anonymous):

\[\LARGE \int\limits \frac{(x+6)dx}{(x+2)^2}\] Use partial fractions

OpenStudy (anonymous):

Write \[ \frac {x+6}{(x+2)^2}= \frac {x+2+4}{(x+2)^2}= \frac {x+2}{(x+2)^2}+ \frac {4}{(x+2)^2}\\ \frac {1}{(x+2)}+ \frac {4}{(x+2)^2} \]

OpenStudy (anonymous):